Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

two maths problems

Expert replies
by real_arsen_ac » Wed Jul 23, 2008 6:54 am
Hey guys, i got 2 problems here in OG:

1.The probability is 1/2 that a certain coin will turn up heads on any given toss. If the coin is to tossed 3 times, what is the probability that on at least one of the tosses the coin will turn up tails?
a.1/8 b.1/2 c.3/4 d.7/8 e.15/16
the answer is d

2.In a nationwide poll, N people were interviewed. If 1/4 of them answered yes to question 1, and of those, 1/3 answered yes to question 2, which of the following expressions represnets the number of people interviewed who did not answer yes to both questions?
a.N/7 b.6N/7 c.5N/12 d.7N/12 e.11N/12
the answer is e
Join the discussion
Source: — Problem Solving |

by Canman » Wed Jul 23, 2008 7:28 am
Q1. You can solve this quickly by calculating the probability that heads will come up 3 times and understanding that P(E) = 1-P(not E).

So P(HHH) = (1/2)*(1/2)*(1/2) = 1/8
and P(not HHH) = 1-(1/8) or 7/8
Join the discussion

by Canman » Wed Jul 23, 2008 7:29 am
I don't know where the smiley came from in my post

but last line should be 1 - 1/8 or 7/8
Join the discussion

by CITI29 » Wed Jul 23, 2008 9:09 pm
Ans for ques 2 shld be 'e'...is it?
Join the discussion

by nitingupta5 » Thu Jul 24, 2008 1:26 am
For the second question it should be

No of Yes for question 1 = N/4
No of Yes for question 2 = N/12(these are the people who said yes to question 1 also)

So there are supposed to be only N-N/12 people who said no to the question

Answer E
Join the discussion